OpenStudy (anonymous):

3x – 2y = 11 4x + 3y = 9 solve using the addition method

OpenStudy (blockcolder):

First of all, either the y's or the x's should have the same coefficient. How can we make this happen?

OpenStudy (anonymous):

-4) 3x – 2y = 11 3) 4x + 3y = 9 -12 x + 8 y = -44 12 x + 9y =27 17 y = -17 y =-1 Replace to get x=3